OpenStudy (anonymous):

cos13+sin13/cos13-sin13 = tanA find A=?

OpenStudy (lgbasallote):

two equal signs?

OpenStudy (hba):

WELCOME TO OPEN STUDY

OpenStudy (lgbasallote):

\[\frac{\cos 13 + \sin 13}{\cos 13 - \sin 13} = \tan A?\]

OpenStudy (anonymous):

I think there is one equal sign.. next sign is for finding..

OpenStudy (anonymous):

If we are just asked to find A then: \[A = \tan^{-1} (\frac{\cos13 + \sin13}{\cos13 - \sin13})\]

OpenStudy (lgbasallote):

i think it's using trig identities...

OpenStudy (lgbasallote):

otherwise that would be too easy...just a job for a calculator lol

OpenStudy (anonymous):

without the use of calculator

OpenStudy (anonymous):

\[\frac{tanA + tanB}{1 - tanA .tanB} = \tan(A + B)\]

OpenStudy (anonymous):

\[A = \tan^{-1} (\frac{1 + \tan13}{1 - \tan13})\] \[A = \tan^{-1} (\frac{\tan45 + \tan13}{1 - \tan45.\tan13}) \implies \tan^{-1}(\tan(45 + 13)) \implies A = 58\]

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

\[\huge \color{green}{\textbf{Welcome To Openstudy..}}\]